### v1

# From Learning to Safety: A Direct Data-Driven Framework for Constrained Control

Abstract
Ensuring safety in the sense of constraint satisfaction for learning-based control is a critical challenge, especially in the model-free case. While safety filters address this challenge in the model-based setting by modifying unsafe control inputs, they typically rely on predictive models derived from physics or data. This reliance limits their applicability for advanced model-free learning control methods. To address this gap, we propose a new optimization-based control framework that determines safe control inputs directly from data. The benefit of the framework is that it can be updated through arbitrary model-free learning algorithms to pursue optimal performance. As a key component, the concept of direct data-driven safety filters (3DSF) is first proposed. The framework employs a novel safety certificate, called the state-action control barrier function (SACBF). We present three different schemes to learn the SACBF. Furthermore, based on input-to-state safety analysis, we present the error-to-state safety analysis framework, which provides formal guarantees on safety and recursive feasibility even in the presence of learning inaccuracies. The proposed control framework bridges the gap between model-free learning-based control and constrained control, by decoupling performance optimization from safety enforcement. Simulations on vehicle control illustrate the superior performance regarding constraint satisfaction and task achievement compared to model-based methods and reward shaping.
{IEEEkeywords}
Learning-based control, safe reinforcement learning, safety filters, control barrier functions.
1 Introduction
1.1 Background
Learning-based control has achieved state-of-the-art performance in addressing complex problems in the presence of uncertainty, including applications in transportation systems [1] and robotics [2]. However, ensuring safety is still a challenging problem, particularly when an explicit model of the system is unavailable. Traditional model-based approaches to safety-critical control, such as model predictive control (MPC) [3], struggle with online computational efficiency and rely on the model, while emerging data-driven methods often lack well-understood safety guarantees.
In control problems, safety is typically defined as maintaining state and input variables within given constraints throughout the system’s evolution. The difficulty lies in the fact that unsafe control policies do not necessarily immediately violate constraints but will lead to constraint violation in the future. As a fundamental principle in safe control, control invariance ensures that a system remains within a safe operating set, contained in the state constraint set. This property is crucial for guaranteeing long-term safety.
A widely adopted approach to enforcing long-term safety is the use of safety filters, which provide a modular framework that can be applied to any control policy, even those without explicit safety considerations [4, 5, 6, 7, 8, 9, 10]. The basic principle of safety filters is to post-process the input of a given control policy such that the resulting closed-loop system remains forward invariant w.r.t. the specified state and input constraints. The design of safety filters typically consists of two phases: an offline phase, where a safety certificate characterizing safe states is computed, and an online phase, where this certificate is incorporated as a constraint to modify potentially unsafe control actions from the reference controller. With the development of different kinds safety certificates, various kinds of safety filters have been proposed, such as control barrier function (CBF)-based safety filters [4, 5, 6], invariant set-based safety filters [7], Hamilton-Jacobi reachability-based safety filters [8, 9], and predictive safety filters [10]. Learning-based approaches have increasingly been used to synthesize safety certificates, particularly for complex, nonlinear systems with non-convex state and input constraints. For a detailed overview, we refer the reader to relevant work in the literature [11, 5, 12, 13, 14] as well as comprehensive surveys [15, 16].
Despite the differences and connections, almost all safety filters rely on a mathematical model, which is either exactly derived from physical principles or approximately estimated. In most formulations, model information is required in both the offline and online phases. Specifically, enforcing invariance conditions, first in the safety certificate, and then in the control policy, requires explicit knowledge of the system dynamics. Recently, there has been a growing number of approaches focusing on offline construction of safety certificates using only state transition data [17]. However, these approaches cannot abandon the reliance on an explicit model during the online phase. This limitation is mainly due to the inherent property of existing safety certificates, which fully work on the state space. In particular, when using an existing safety certificate, safety filters need a prediction model to enforce safety conditions on the successor states. A detailed explanation is provided at the end of Section II.
To overcome the above limitation, data-driven safety filters have received significant attention. Almost all existing data-driven safety filters belong to indirect data-driven methods [18, 15, 8, 13, 14, 19, 10]. The difference between direct and indirect methods was made in the context of adaptive control community. In indirect approaches, first system identification or disturbance estimation is performed and then a controller is learned based on the obtained model. In direct methods the controller is learned directly from data, bypassing the need for system identification. Indirect data-driven safety filters have one main issue: The errors arising from both model identification and certificate learning will compound, leading to a degradation in the safety performance of the filtered controller. Among all data-driven safety filters, there is only one direct data-driven formulation, which learns discriminating hyper-planes to directly regulate the control inputs [20]. However, this method is limited to linear safety constraints on inputs, potentially leading to conservative control actions if nonlinear constraints are considered, and lacks formal guarantees regarding constraint satisfaction.
An alternative approach bypasses model identification by jointly learning a CBF and an explicit policy that enforces the CBF constraint, but this often results in overly conservative policies focused solely on safety [21, 16].
Similar to the distinction between indirect and direct data-driven control, learning-based control can also be categorized into model-based learning and model-free learning. Learning-based control, encompassing supervised learning (SL) and reinforcement learning (RL) [22], iteratively finds an optimal control policy that minimizes a pre-defined cost. Due to the stochastic nature of learning algorithms, learning-based control, especially in the absence of an explicit model, cannot fully guarantee safety without using safety filters to regulate policy execution. However, as almost all data-driven safety filters still rely on an underlying model, there still remains a gap when applying data-driven safety filters to learning-based control approaches that does not use an explicit model.
(a) The flow chart of indirect data-driven safe control using safety filters involving classical safety certificates.
(b) The flow chart of the proposed 3DSF, a special form of the optimization-based control framework we propose.
Figure 1: The comparison between the proposed 3DSF and the existing indirect data-driven counterparts.
In this paper, we focus on designing a direct data-driven safety filter (3DSF) based on our previously proposed concept of state-action control barrier functions (SACBFs) [23]. Unlike classical safety certificates, which only evaluate safety in the state space, the SACBF framework enables safety evaluation for each state-action pair. This key feature eliminates the need for explicit system dynamics during policy modification, making it particularly suitable for model-free learning-based control. The comparison between the proposed 3DSF and the existing indirect data-driven counterparts is illustrated in Fig. 1. For the indirect data-driven methods, both system identification and certificate learning should be performed one by one or simultaneously. The identified model and learned certificate are both used in the safety filter design. In comparison, the proposed 3DSF only uses an SACBF. A qualitative comparison of our method with others can be found in Section VII.
1.2 Contributions
The paper contributes to the state of the art in the following aspects:
(1) Optimization-based direct data-driven safe control: We propose a novel safe control framework for general nonlinear systems with nonlinear constraints. The main advantage is that the safe controller can be trained and implemented using state transition data only, without the need for system identification. The control framework incorporates a novel safety filter, which we call the direct data-driven safety filter (3DSF). This framework, which can be integrated with arbitrary learning-based controller synthesis methods, also provides formal guarantees on constraint satisfaction.
(2) Learning-based synthesis of SACBFs: We develop three distinct learning-based methods for synthesizing SACBFs directly from data: including SL from a known CBF, learning from an expert safe controller, and RL (self-learning).
(3) Robustness analysis via error-to-state safety (ESSf): We propose a systematic framework, called Error-to-State Safety (ESSf), to analyze how learning-induced errors in SACBFs affect overall safety performance. The framework motivates the approach of state constraint tightening followed by SACBF constraint relaxation to ensure that the learned SACBF-regulated controller meets safety requirements.
(4) A unified framework extending RL to constrained control: Building on the proposed optimization-based safe control, we present a general strategy to extend classical unconstrained value-based RL algorithms to constrained ones so that suboptimal performance regarding other tasks is achieved. This reversely shows another benefit of our proposed optimization-based approach: it separates the learning of the controller into two components: optimizing performance through learning the objective function, and ensuring safety through learning the constraints.
Additionally, for contribution (2), if system nonlinearities are known, we formulate the synthesis of a quadratic SACBF as solving a convex optimization problem with linear matrix inequality (LMI) constraints.
2 Preliminaries and problem formulation
2.1 Notations
The set $\mathbb{R}$ ($\mathbb{R}_{\geq 0}$) is the set of (nonnegative) real numbers, and the set of real vectors with the dimension $n$ is denoted by $\mathbb{R}^{n}$. The sets $\mathbb{N}$ and $\mathbb{N}^{+}$ represent the set of non-negative integers and the set of positive integers, respectively. Besides, $\mathbb{N}_{a}=\{0,1,...,a\}$, and $\mathbb{N}^{+}_{a}=\{1,...,a\}$ for any positive integer $a$. The matrix $I_{n}$ is the identity matrix with the dimension $n\times n$. The notation $A^{\dagger}$ represents the right inverse of the matrix $A$. The relation $A\succeq 0$ means that the matrix $A$ is positive semi-definite. The determinant of a matrix $A$ is denoted by $\mathrm{det}(A)$. The dimension of a vector $x$ is denoted by $\mathrm{dim}(x)$. The norm $||x||_{2}$ is the Euclidian norm of the vector $x$. The number of elements in a finite set $S$ is called its cardinality, denoted by $|S|$. The set $\mathcal{B}_{\epsilon}(\bar{z}):=\{z\in\mathbb{R}^{n_{z}}\mid||z-\bar{z}||_{2}\leq\epsilon\}$ denotes the closed ball around $\bar{z}$ with radius $\epsilon$. The unit step function $\mathrm{step}(\cdot)$ returns 1 if the input is larger than 0 and returns 0 otherwise. A continuous function $\alpha(\cdot):[0,a)\to[0,\infty)$ for some $a\in(0,\infty]$ is said to belong to class $\mathcal{K}$ if it is strictly increasing and $\alpha(0)=0$. A class $\mathcal{K}$ function $\alpha$ is said to belong to class $\mathcal{K}_{\infty}$ if it further satisfies $a=\infty$ and $\alpha(r)\to\infty$ as $r\to\infty$.
2.2 Preliminaries
We consider a deterministic discrete-time nonlinear system
$x_{t+1}=f\left(x_{t},u_{t}\right),\quad t=0,1,\ldots,$ (1)
where $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$ are the state and the input at time step $t$, and $f(\cdot,\cdot):\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is a continuous function. We consider a constrained optimal control problem, in which the states and inputs should satisfy time-invariant constraints: $x_{t}\in X:=\{x\in\mathcal{X}|h(x)\leq 0\}$ and $u_{t}\in U\subseteq\mathcal{U}$ for all time steps. Here, $h(\cdot):\mathcal{X}\to\mathbb{R}$ is a scalar-valued continuous function that defines the state constraint [^1] For the constraint defined by multiple inequalities $h_{i}(x)\leq 0,\;i=1,2,...,I$, we can let $h(x)=\max_{i\in\mathbb{N}^{+}_{I}}h_{i}(x)$. The set $\{x|h(x)\leq 0\}$ is identical to $\{x|h_{i}(x)\leq 0,\;\forall i\in\mathbb{N}^{+}_{I}\}$, and $h$ will be continuous if each $h_{i}$ is continuous.. The set $U$ is compact. For the convenience of performance analysis and sampling, we require the compactness of $\mathcal{X}$ and $X$. In our framework, we do not assume the knowledge of the explicit form of the system dynamics $f(\cdot,\cdot)$. Instead, we require the availability of transition samples $(x_{t},u_{t},x_{t+1})$, achieved through simulation or experimental methods. Various sampling strategies, including random sampling and grid-based sampling, may be employed to obtain the transition triples.
Given an initial state $x_{0}$, we are interested in designing a deterministic control policy $\pi:\mathcal{X}\to\mathcal{U}$ to steer the trajectory of the system to a non-empty target region $X_{\mathrm{tar}}\subseteq X$. To achieve the control objective, for any state $x$ and input $u$, a stage cost is defined by $g\left(x,u\right)$, which is non-negative and upper-bounded. For any policy $\pi$, the value function $J^{\pi}\left(\cdot\right):\mathcal{X}\to[0,\infty)$ is defined by
[…]
where $\gamma\in(0,1)$ is a discount factor. The objective is to find an optimal deterministic policy $\pi^{*}(\cdot):\mathcal{X}\to\mathcal{U}$ that solves the following infinite-horizon optimal control problem:
$J^{*}\left(x\right):=\inf_{\pi}J^{\pi}\left(x\right)\;\mathrm{s.t.}\;h(x_{t})\leq 0,\pi(x_{t})\in U,t=0,1,\ldots,$ (3)
where $J^{*}(\cdot):\mathcal{X}\to[0,\infty)$ is the optimal value function.
The exact form of the optimal policy $\pi^{*}$ is difficult to compute due to the following reasons: (i) The number of constraints in (3) is infinite; (ii) The closed-form expressions of the objective function are unknown since the system is unknown and the horizon is infinite. To get an approximation of $\pi^{*}(\cdot)$, a parameterized policy is usually preferred in literature and then the parameters updated using RL or SL [24, 22].
2.3 Control barrier functions for safety
To ensure that the learned policy satisfy the constraints in (3), two different classes of methods, including cost shaping [25] and using safety certificates [26, 5, 4], have been developed in recent papers. In contrast to the cost shaping method, which usually does not provide strict safety guarantees, using safety certificates can impose a constraint on the control input to ensure safety. The CBF is one of the most popular safety certificates.
Definition 1 (Control barrier function [26]).
A function $B(\cdot):\mathcal{X}\to\mathbb{R}$ is called a control barrier function (CBF) with a corresponding safe set $\mathcal{S}_{B}:=\left\{x\in\mathbb{R}^{n_{x}}|B(x)\leq 0\right\}\subseteq\mathcal{X}$, if $\mathcal{S}_{B}$ is non-empty, $h(x)\leq B(x),\;\forall x\in\mathcal{X}$, and if there exists a $\beta_{B}\in[0,1]$ such that
$\min_{u\in U}B(f(x,u))\leq\beta_{B}B(x),\;\forall x\in\mathcal{X}.$ (4)
With a CBF available, one can implicitly construct a policy $\pi(\cdot)$ as an optimizer of the following nonlinear optimization problem:
$\pi(x):=\arg\min_{u\in U}$ $\;Q(x,u)$
$\mathrm{s.t.}\;$ $B(f(x,u))\leq\beta_{B}B(x),$ (5)
where $Q:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is a problem-dependent objective function. For instance, one can specify $Q$ as $||u-\pi^{\mathrm{unsafe}}(x)||_{2}$ where $\pi^{\mathrm{unsafe}}$ is a reference policy that may have some other acceptable performance. With this specification, (2.3) works as a safety filter to refine the potentially unsafe policy $\pi^{\mathrm{unsafe}}$. One can also let $Q$ be an approximation of the state-action optimal value function (Q function) of (3), that is,
$Q^{*}(x,u):=g(x,u)+\gamma J^{*}(f(x,u)),$ (6)
which is commonly used in RL. In this situation, $\pi$ becomes the greedy policy w.r.t. the optimal value function under a given CBF constraint.
A valid CBF is sufficient to guarantee the safety of $\pi$ [4]. However, the main limitation of (2.3) is that the knowledge of $f$ is required. Even though some data-driven methods exist for learning CBFs from black-box models [17], system identification is necessary to implement (2.3). This limitation significantly restricts the application of safe filters and CBFs in reinforcement learning algorithms that do not require an explicit model. Moreover, the mismatch between the identified model and the real system, along with inaccuracies in learning CBFs, jointly affect the safety performance of $\pi$.
3 State-action CBFs and control parameterization
Inspired by the Q function in RL, we propose a new optimization-based control framework that does not contain $f$:
$\pi(x):=\arg\min_{u\in U}$ $\;Q(x,u)$
$\mathrm{s.t.}\;$ $Q^{B}(x,u)\leq 0,$ (7)
where $Q$ can be designed similarly to that in (2.3), and another function $Q^{B}(\cdot,\cdot):\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is used to add a constraint to regulate the behavior of $\pi$. To make $\pi$ satisfy the safety requirements $h(x_{t})\leq 0,\;\forall t\in\mathbb{N}$, we consider the definition of state-action CBFs, which we have introduced in [23].
Definition 2 (State-action control barrier function (SACBF)).
A function $Q^{B}(\cdot,\cdot):\mathcal{X}\times U\to\mathbb{R}$ is called a state-action control barrier function (SACBF) with a corresponding safe set $\mathcal{S}_{Q}$ of states, if the pair $(Q^{B},\;\mathcal{S}_{Q})$ satisfies the following conditions:
(i) $\mathcal{S}_{Q}$ is non-empty, and $h(x)\leq 0,\;\forall x\in\mathcal{S}_{Q}$.
(ii) $\min_{u\in U}Q^{B}(x,u)\leq 0,\;\forall x\in\mathcal{S}_{Q}$.
(iii) For any $x\in\mathcal{S}_{Q}$, any $u\in U$ satisfying $Q^{B}(x,u)\leq 0$ ensures that $f(x,u)\in\mathcal{S}_{Q}$.
There are two main features that distinguish SACBFs from standard CBFs. First, the SACBF is a function of both the states and the input. It not only characterizes the energy of a state but also quantifies the quality (safety) of selecting an input in a given state. Besides, the explicit form of the safe set is not specified in Definition 2.
According to the definition of SACBFs, we have the following properties for (3):
Lemma 1 (Safety of SACBFs[23]).
Considering the policy $\pi$ in (3), if $Q^{B}$ is an SACBF with the safe set $\mathcal{S}_{Q}$, $\pi$ will render (1) positively invariant in $\mathcal{S}_{Q}$. As a result, the trajectories of (1), starting from $x_{0}\in\mathcal{S}_{Q}$, controlled by $\pi$, satisfy $h(x_{t})\leq 0$ and $u_{t}\in U,\;\forall t\in\mathbb{N}$.
Remark 1.
Analogous to CBFs, SACBFs are difficult to compute exactly for nonlinear problems. In the remainder of this paper, we will explore using learning-based methods to synthesize the optimization-based controller, i.e., to learn the functions $Q$ and $Q^{B}$. A tractable way is to parameterize them by $Q_{\theta}$ and $Q^{B}_{\omega}$ with the parameters $\theta$ and $\omega$, and then learn these parameters. Instead of simultaneously updating $Q$ and $Q^{B}$, we prefer to first learn an SACBF $Q^{B}_{\omega}$ (Section IV) and subsequently to integrate it into the learning of the objective function $Q_{\theta}$ to achieve the optimal control objective (Section VI). This strategy is motivated by the observation that safety is inherently independent of other performance metrics, whereas achieving optimality regarding task performance should be addressed under the premise of ensuring safety.
The policy $\pi$ in (3) with the parameterization $Q_{\theta}$ and $Q^{B}_{\omega}$ is denoted by $\pi_{\theta,\omega}$.
4 Learning state-action control barrier functions
In this section, we propose three learning-based approaches to approximate an SACBF $Q^{B}_{\omega}$. For the first one, we assume the knowledge of a valid CBF and use SL to obtain an SACBF. The second approach, inspired by [11], employs sampling-based methods to approximate the solution of a robust optimization problem. Such a solution is guaranteed to lead to a valid SACBF. This approach requires prior knowledge of a safety control policy (usually conservative w.r.t. task performance). For the third approach, we connect the synthesis of SACBFs with Hamilton-Jacobi reachability [27], and thereby propose an RL-based approach to obtain $Q^{B}_{\omega}$. The third approach only relies on the availability of state transition data. Besides, when the nonlinearity of the system is known, we also find that the synthesis of SACBFs can be achieved by solving a convex optimization problem with LMI conditions.
4.1 Supervised learning SACBFs from CBFs
We first present our simplest learning-based method for obtaining an SACBF, based on a rather restrictive assumption that we have a CBF at hand. In certain applications such as adaptive cruise control [4], CBFs can be manually designed based on state constraints, such as the distance between adjacent vehicles. However, as noted in the introduction, such CBFs cannot be used to design a safety filter (2.3) when the model is not fully known. In this subsection, we explore connections between SACBFs and CBFs and demonstrate that one can easily learn an SACBF from a given CBF.
Assumption 1.
A CBF $B(\cdot):\mathcal{X}\to\mathbb{R}$ is known.
Proposition 1.
Under the constraints $h(x)\leq 0$ and $u\in U$, suppose that $B$ is a CBF satisfying (4) with $\beta_{B}\in[0,1]$, and that the safe set is $\mathcal{S}_{B}$. Then, $Q^{B}$ defined by
$Q^{B}(x,u):=\max\{h(x),\;\frac{1}{\beta_{B}}B(f(x,u))\}$ (8)
is an SACBF with the safe set $\mathcal{S}_{Q}=\mathcal{S}_{B}$.
Proof.
From the relation between $Q^{B}$ and $B$, we can get that $h(x)\leq\min_{u\in U}Q^{B}(x,u)\leq 0,\;\forall x\in\mathcal{S}_{B}$. The above arguments prove that $Q^{B}$ obeys (i) and (ii) of Definition 2. Furthermore, for any $x\in\mathcal{S}_{B}$, if $Q^{B}(x,u)\leq 0$, we have $B(f(x,u))\leq 0$ and thus $f(x,u)\in\mathcal{S}_{B}$. This finishes the proof. ∎
Proposition 1 implies a straightforward way to learn an SACBF. Initially, a collection of transition triples $\{(x^{(i)},u^{(i)},f(x^{(i)},u^{(i)}))\}^{N}_{i=1}$, is generated from the state space $\mathcal{X}$ and the control space $U$. Subsequently, the labels $Q^{B}(x^{(i)},u^{(i)})$ can be computed utilizing (8). Following this, a regression model $Q^{B}_{\omega}$ is constructed and trained to minimize the empirical loss, specifically the mean squared error between $Q^{B}_{\omega}$ and the computed labels.
4.2 Learning SACBFs from an expert controller
In the presence of both state and input constraints, manually crafted CBFs often fail. Towards the goal of learning a valid SACBF, like in [11] where CBFs are synthesized from expert demonstrations, we assume the availability of an expert safe controller $\pi_{\mathrm{s}}(\cdot):\mathcal{X}\to\mathcal{U}$ in this subsection. This assumption is reasonable in certain scenarios, such as human-controlled systems (e.g., driving), where safety is achieved implicitly without an explicit model.
Formally, we have the following assumption in this subsection.
Assumption 2.
There exists a continuous policy $\pi_{\mathrm{s}}$ and a compact set $\mathcal{S}_{0}\subseteq X$ ($\mathcal{S}_{0}$ can be unknown) such that with the initial condition $x_{0}\in\mathcal{S}_{0}$, the state-input trajectories of the system (1) with $\pi_{\mathrm{s}}$ always stay in $\mathcal{S}_{0}\times U$ and such that the states reach $X_{\mathrm{tar}}$ in a finite number of time steps $T$.
Under Assumption 2, we formulate the synthesis of $Q^{B}$ as the following optimization problem:
[…]
s.t. $0\leq q(x),\;\forall x\in\mathcal{X}\setminus\mathcal{S}_{0}$ (9b)
$Q^{B}(x,\pi_{\mathrm{s}}(x))\leq\beta q(x),\forall x\in\mathcal{S}_{0}$ (9c)
$q(f(x,u))\leq Q^{B}(x,u),\forall(x,u)\in\mathcal{S}_{0}\times U$ (9d)
where $Q^{B}$ and $q(\cdot):\mathcal{X}\to\mathbb{R}$ are continuous in their domain, and $\beta\in[0,1)$ is a tuning parameter.
Proposition 2 (Converse SACBFs).
Under Assumption 2, there exists a $\beta\in[0,1)$ such that the problem (9) is feasible and such that any optimal solution pair $(Q^{B*},q^{*})$ ensures that $Q^{B*}$ is an SACBF with the safe set $\mathcal{S}^{*}_{Q}=\{x\in\mathcal{X}|q^{*}(x)\leq 0\}$.
Proof.
The proof consists of two parts. In the first part, we will prove the feasibility of (9). By fixing $Q(x,u)=q(f(x,u))$, which satisfies (9d), the constraint (9c) becomes $q(f(x,\pi_{\mathrm{s}}(x)))\leq\beta q(x),\;\forall x\in\mathcal{S}_{0}$.
Since $\mathcal{S}_{0}$ is compact, there exists a continuous function $q$ such that (9b) and
$\mathcal{S}_{0}$ =$=\left\{x\in\mathcal{X}|q(x)\leq 0\right\},$
$\partial\mathcal{S}_{0}$ =$=\left\{x\in\mathcal{X}|q(x)=0\right\},$
$\operatorname{Int}(\mathcal{S}_{0})$ =$=\left\{x\in\mathcal{X}|q(x)<0\right\}$ (10)
hold. Actually, the above statement can be proved by considering $q(x)=\max\{h(x),\;\bar{q}(x)\}$, where $\bar{q}$ is a distance function defined by
$\bar{q}(x)=\begin{cases}-\inf_{y\in\partial\mathcal{S}_{0}}\|x-y\|\text{ if }x\in\mathcal{S}_{0}\\ \inf_{y\in\partial\mathcal{S}_{0}}\|x-y\|\text{ otherwise}.\end{cases}$ (11)
With the above definitions, we follow the proof of the Converse CBF Theorem in the continuous time domain [4, Proposition 3] to prove the feasibility of (9). In particular, for any continuous $q$ such that (9b) and (4.2) hold, define a function $\alpha(\cdot):[0,\infty)\to\mathbb{R}$ by
$\alpha(r)=\sup_{\{x\in\mathcal{X}|-r\leq q(x)\leq 0\}}q(f(x,\pi_{\mathrm{s}}(x)))-q(x).$ (12)
Since $\{x\in\mathcal{X}|-r\leq q(x)\leq 0\}$ is compact and $q$, $f$, and $\pi_{\mathrm{s}}$ are continuous, $\alpha$ is well-defined and non-decreasing, and satisfies $q(f(x,\pi_{\mathrm{s}}(x)))-q(x)\leq\alpha(-q(x)),\;\forall x\in\mathcal{S}_{0}$.
When $x$ is such that $q(x)=0$, i.e., $x\in\partial\mathcal{S}_{0}$, by the invariance of $\mathcal{S}_{0}$, we have $q(f(x,\pi_{\mathrm{s}}(x)))\leq 0$, and consequently $\alpha(0)\leq 0$. Meanwhile, by taking the extreme values 00 and $-r$ of $q(f(x,\pi_{\mathrm{s}}(x)))$ and $q(x)$ respectively, we have $\min_{r>0}\frac{\alpha(r)}{r}\leq\min_{r>0}\frac{0-(-r)}{r}=1$, which, combined with $\alpha(0)\leq 0$, implies that there exists a class $\mathcal{K}$ function $\bar{\alpha}$ upper-bounding $\alpha$ and satisfying
$q(f(x,\pi_{\mathrm{s}}(x)))-q(x)\leq\bar{\alpha}(-q(x)),\forall x\in\mathcal{S}_{0},$
$\min_{r>0}\frac{\bar{\alpha}(r)}{r}\leq 1.$ (13)
Letting
$\beta=1-\min_{r>0}\frac{\bar{\alpha}(r)}{r}\in[0,1)$ (14)
proves the feasibility of (9).
In the second part of the proof, we will show that the optimal solution to (9) yields an SACBF. To this end, we first establish the non-emptiness of $\mathcal{S}^{*}_{Q}$ by contradiction. Suppose that $\mathcal{S}^{*}_{Q}$ is empty, i.e., $q^{*}(x)>0,\;\forall x\in\mathcal{X}$. Consider a new function $q^{\prime}(\cdot):\mathcal{X}\to\mathbb{R}$ defined by $q^{\prime}(x)=\min\left\{q^{*}(x),\;\max\{h(x),\;\bar{q}(x)\}\right\}$, where $\bar{q}$ is defined by (11). Based on the proof of the feasibility of (9), $\max\{h(x),\;\bar{q}(x)\}$ satisfies all the constraints in (9). Therefore, we have
$q^{\prime}(x^{+})\!\leq\!\min\left\{\beta q^{*}(x),\beta\max\{h(x),\bar{q}(x)\}\right\}\!\leq\!\beta q^{\prime}(x),\;\forall x\in\mathcal{S}_{0}$
[…]
The condition (i) of Definition 2 holds because of (9b) and the non-emptiness of $\mathcal{S}^{*}_{Q}$. Moreover, $\forall x\in\mathcal{S}^{*}_{Q}$, it follows from (9c) that $\min_{u\in U}Q^{B*}(x,u)\leq Q^{B*}(x,\pi_{\mathrm{s}}(x))\leq\beta q(x)\leq 0$, meaning that the condition (ii) holds. Furthermore, (9d) enforces the condition (iii). These complete the proof that $Q^{B*}$ is an SACBF. ∎
For computational tractability, $Q^{B}$ and $q$ in (9) are parameterized as $Q^{B}_{\omega}$ and $q_{\omega}$, with all the parameters condensed in $\omega$. As a result, the search space becomes the parameter space of $\omega$. To solve (9), we use state transition samples to relax the constraints that should hold in the continuous state(-input) space to constraints that only need to hold for the samples.
In particular, let $\{x^{(i)}_{0}\}^{N}_{i=1}$ denote the set of initial states. Here, $N$ is the number of samples. For each $x^{(i)}_{0}$, we apply $\pi_{\mathrm{s}}$ to (1) for $T$ time steps and get the state-input trajectory $\left\{(x^{(i)}_{t},\pi_{\mathrm{s}}(x^{(i)}_{t}))\right\}^{T}_{t=0}$. If this trajectory obeys $x^{(i)}_{t}\in X,\;\pi_{\mathrm{s}}(x^{(i)}_{t})\in U,\;\forall t\in\mathbb{N}_{T}$, and $x^{(i)}_{T}\in X_{\mathrm{tar}}$, then $x^{(i)}_{t}\in\mathcal{S}_{0},\;\forall t\in\mathbb{N}_{T}$. Otherwise, $x^{(i)}_{t}\notin\mathcal{S}_{0},\;\forall t\in\mathbb{N}_{T}$. The above checking procedure allows us to separate the state trajectories according to whether or not they are in $\mathcal{S}_{0}$, without knowing $\mathcal{S}_{0}$. Then, we define the set $\mathcal{S}_{\mathrm{s}}:=\{x^{(i)}_{t}|x^{(i)}_{t}\in X,\;\pi_{\mathrm{s}}(x^{(i)}_{t})\in U,\;\forall t\in\mathbb{N}_{T}$, $x^{(i)}_{T}\in X_{\mathrm{tar}},\;\forall i\in\mathbb{N}^{+}_{N}\}$, which contains the safe state trajectories. Meanwhile, we define $\mathcal{X}_{\mathrm{s}}:=\{x^{(i)}_{t}\}^{N\;\;\;T}_{i=1\;t=0}$.
For the optimization problem (9), we replace the infinite sets $\mathcal{S}_{0}$, $\mathcal{X}$, and $\mathcal{S}_{0}\times U$ by the finite sets $\mathcal{S}_{\mathrm{s}}$, $\mathcal{X}_{\mathrm{s}}$, and $\mathcal{S}_{\mathrm{s}}\times U_{\mathrm{s}}$, respectively. Here $U_{\mathrm{s}}$ is the set of inputs sampled in $U$ through some sampling strategy. As a result, (9) is simplified to a constrained optimization problem with a finite number of constraints. If $Q^{B}_{\omega}$ and $q$ are linear in $\omega$, the problem (9) becomes linear and can be solved efficiently. In a more general case such as when these functions are represented by deep neural networks, a tractable solution involves transforming the constraints into soft penalties and adding them into the objective.
Remark 2.
Due to the above constraint relaxation, the validity of the approximated solution is however not guaranteed. To get a valid SACBF, we can follow the approach of [11] to tighten the constraints in (9). Assuming $Q^{B}_{\omega}$, $q_{\omega}$, $\pi_{\mathrm{s}}$, and $f$ to be Lipschitz as well as having sufficient sampling, $Q^{B}_{\omega}$ will be a valid SACBF222“Sufficient sampling” means that $\mathcal{X}_{\mathrm{s}}\times U_{\mathrm{s}}$ constitutes an $\epsilon$-net of $\mathcal{X}\times U$ with $\epsilon>0$ [11]. Namely, $\forall x\in\mathcal{X}$ and $\forall u\in U$, $\exists x^{\prime}\in\mathcal{X}_{\mathrm{s}}$ and $\exists u^{\prime}\in U_{\mathrm{s}}$ such that $||x-x^{\prime}||^{2}_{2}+||u-u^{\prime}||^{2}_{2}\leq\epsilon^{2}$.. Although this constraint tightening approach will guarantee that $Q^{B}_{\omega^{*}}$ is a valid SACBF, it may be conservative under insufficient sampling. Furthermore, estimating the Lipschitz constants is a non-trivial task. To address this limitation, we will propose a new tightening approach exploits the inherent robustness of the SACBF. This will be elaborated on in Section 5.
4.3 Learning SACBFs via RL
In the previous subsection, we have proposed an optimization-based approach to learn SACBFs from a known safe controller. This approach has two main limitations. First, a safe controller is sometimes hard to obtain as a prior. Second, the safe set of the learned SACBF cannot be larger than the safe region $\mathcal{S}_{0}$ of the safe controller. On the other hand, it is desirable to learn an SACBF with the largest possible safe set directly from randomly generated transition data, without any expert supervision. For this purpose, we observe that the proposed SACBFs show some similarities with state-action value functions (Q functions) in RL. These similarities motivate us to use RL methods to synthesize SACBFs.
We consider the optimal value function of the following optimal control problem:
$B^{*}(x):=$ $\min_{\{(x_{t},u_{t})\}_{t=0}^{\infty}}\max_{t\in\mathbb{N}}h(x_{t})$
$\quad\mathrm{s.t.}\;(1),\;\;u_{t}\in U,\;t\in\mathbb{N},\text{ and }x_{0}=x,$ (15)
which is a typical reachability problem [27] in the discrete-time domain.
Remark 3.
The optimal value function $B^{*}$, which can be obtained by dynamic programming, is a CBF with the safe set being the maximal control-invariant set [9]. Using Proposition 1 and letting $\beta_{B}=1$, we know that the state-action optimal value function of (4.3), defined by
$Q^{B*}(x,u):=\max\{h(x),\;B^{*}(f(x,u))\},$ (16)
is an SACBF, of which the safe set is also the maximal control-invariant set.
Combining (4.3) and (16), it is observed that $Q^{B*}$ is one solution of the Bellman optimality equation [22]:
$Q^{B}(x,u)=\max\{h(x),\min_{u^{+}\in U}Q^{B}(f(x,u),u^{+})\}.$ (17)
For general nonlinear systems, accurately computing $Q^{B*}$ from (4.3) and (16) is nearly impossible. Similar to the proposed SL method and the proposed expert-guided learning method, an approximation of $Q^{B*}$ is necessary. Since the form of $f$ is unknown, it is not possible to design query-based algorithms and use SL methods to get the approximation. However, by employing (17) and the theoretical results in Proposition 1, we can adapt off-the-shelf value-based RL methods such as temporal difference learning methods [28, 29] and neural fitted Q iteration (FQI) [30] to approximate $Q^{B*}$.
Before learning $Q^{B*}$, it is crucial to recognize that (17) may have multiple solutions and that most RL methods based on Bellman iteration could approach any solution of (17). The following result eliminates our concern by stating that any solution of (17), satisfying $\min_{x\in\mathcal{X},u\in U}Q^{B}(x,u)\leq 0$, is an SACBF.
Proposition 3.
Consider the Bellman optimality equation (17). Any solution $Q^{B}$ satisfying $\min_{x\in\mathcal{X},u\in U}Q^{B}(x,u)\leq 0$ is an SACBF with the safe set $\mathcal{S}_{Q}=\{x\in\mathbb{R}^{n_{x}}|\min_{u\in U}Q^{B}(x,u)\leq 0\}$.
Proof.
It follows from $\min_{x\in\mathcal{X},u\in U}Q^{B}(x,u)\leq 0$ that $\mathcal{S}_{Q}$ is non-empty. From (17), we deduce that $h(x)\leq\min_{u\in U}Q^{B}(x,u)\leq 0,\;\forall x\in\mathcal{S}_{Q}$. The above arguments prove that $Q^{B}$ satisfies (i) of Definition 2. The second condition of Definition 2 directly follows from $\mathcal{S}_{Q}=\{x\in\mathbb{R}^{n_{x}}|\min_{u\in U}Q^{B}(x,u)\leq 0\}$. Finally, if $Q(x,u)\leq 0$, by (17) we have $\min_{u^{+}\in U}Q^{B}(f(x,u),u^{+})\leq 0$, which means that the successor state $f(x,u)\in\mathcal{S}_{Q}$. ∎
Following the temporal difference learning method in [28], the training loss for the candidate SACBF $Q^{B}_{\omega}$ is designed as
$l(\mathcal{D},\omega)$ $=l_{1}(\mathcal{D},\omega)+\rho l_{2}(\mathcal{D},\omega),\;\mathrm{with}$
$l_{1}(\mathcal{D},\omega)$ $=\sum_{(x,u,x^{+})\in\mathcal{D}}\left.(\max\{h(x),\min_{u^{+}\in U}Q^{B}_{\omega}(x^{+},u^{+})\}\right.$
$\quad\quad\quad\quad\quad\quad\left.-Q^{B}_{\omega}(x,u)\right)^{2}$
$l_{2}(\mathcal{D},\omega)$ $=\sum_{(x,u,x^{+})\in\mathcal{D}}Q^{B}_{\omega}(x,u),$ (18)
where $\mathcal{D}=\{(x^{(i)},u^{(i)},f(x^{(i)},u^{(i)}))\}^{N}_{i=1}$ represents the collection of state transition triples. Intuitively, minimizing the loss $l_{1}$ encourages $Q^{B}_{\omega}$ to reduce the temporal difference, thereby approaching the solution of (17). Meanwhile, as the Bellman equation (3) may have multiple solutions, the loss $l_{2}$ is introduced to guide $Q^{B}_{\omega}$ toward the smallest solution of (17), which has the largest safe set. The positive parameter $\rho$ balances the relative importance of the two loss functions.
4.4 Special case when nonlinearity is known
The methods for synthesizing SACBFs proposed above are aimed at general nonlinear systems. When the nonlinearity of the system is known, we can leverage formulas from data-driven control [31, 32] and compute SACBFs through convex optimization. Formally, in this subsection, we require the system to satisfy the following conditions:
Assumption 3.
  • The system is described by the following equation:
    $x_{t+1}=A\bar{v}(x_{t})+Bu_{t},$ (19)
    where $\bar{v}(x)=\left[\begin{array}[]{c}x\\ v(x)\end{array}\right]$ includes a known nonlinear function $v(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{v}}$. Besides, the linear system obtained by linearizing (19) at the origin is controllable in the absence of the constraints.
  • The constraints are linear, i.e., $X=\{x\in\mathbb{R}^{n_{x}}|H_{x}x\leq h_{x}\},\;U=\{u\in\mathbb{R}^{n_{u}}|H_{u}u\leq h_{u}\}$. Here, $H_{x}$ and $H_{u}$ are matrices with appropriate dimensions. Besides, $h_{x}$ and $h_{u}$ are two vectors with strictly positive elements. This indicates that the origin is contained in the interior of $X\times U$.
  • The target region $X_{\mathrm{tar}}=\{0\}$.
To begin with, we recall the definition of Hankel matrix [31] associated to any time-varying signal $\{z_{t}\}_{t\in\mathbb{N}}$ with $z_{t}\in\mathbb{R}^{n_{z}},\;t\in\mathbb{N}$:
$\mathbf{z}_{i,j,k}=\left[\begin{array}[]{cccc}z_{i}&z_{i+1}&\cdots&z_{i+k-j}\\ z_{i+1}&z_{i+2}&\cdots&z_{i+k-j+1}\\ \vdots&\vdots&\ddots&\vdots\\ z_{j+i-1}&z_{j-i}&\cdots&z_{k+i-1}\end{array}\right].$ (20)
Then, the Hankel matrices $\mathbf{x}_{i,j,k}$ and $\mathbf{u}_{i,j,k}$ contain the state and input data collected from a simulation of the system in (19). The following lemma shows that (19) has equivalent data-based representations.
Lemma 2 ([32]).
Let $\mathbf{D}:=\left[\begin{array}[]{l}\mathbf{u}_{0,1,k}\\ {\mathbf{\bar{v}}_{0,1,k}}\end{array}\right]=\left[\begin{array}[]{l}\mathbf{u}_{0,1,k}\\ \mathbf{x}_{0,1,k}\\ \mathbf{v}_{0,1,k}\end{array}\right]$. Suppose that $\mathrm{rank}(\mathbf{D})=n_{x}+n_{u}+n_{v}$ and that Assumption 3 holds. Then, the system in (19) has the following equivalent representation:
$x_{t+1}=\mathbf{x}_{1,1,k}\mathbf{D}^{\dagger}\left[\begin{array}[]{l}u_{t}\\ \bar{v}(x_{t})\end{array}\right].$ (21)
Furthermore, the system in closed loop with $u=K\bar{v}(x)$ has the following equivalent representation:
$x_{t+1}=\mathbf{x}_{1,1,k}G\bar{v}(x_{t})$ (22)
with $G=[G_{1}\;G_{2}]$, $G_{1}\in\mathbb{R}^{k\times n_{x}}$, and $G_{2}\in\mathbb{R}^{k\times n_{v}}$satisfying
$\left[\begin{array}[]{l}K\\ I_{n_{x}+n_{v}}\end{array}\right]=\mathbf{D}G.$ (23)
In Lemma 2, the rank condition on $\mathbf{D}$ is referred to as the condition of persistency of excitation, which is commonly assumed among direct data-driven control approaches [31]. This condition implies that the collected data is sufficiently informative to represent the system behavior. Next, we show that under persistency of excitation and a mild assumption on the nonlinear term $v$ (See Assumption 4 below), we can formulate a data-driven computation of an SACBF in terms of a convex optimization problem subject to LMI constraints.
Assumption 4.
There exist a constant $\eta>0$ and a matrix $M\in\mathbb{R}^{n_{x}\times n_{x}}$ such that $v^{T}(x)v(x)\leq\eta x^{T}M^{T}Mx$ $\forall x\in X$.
This assumption limits the magnitude of the nonlinear term $v(x)$ by a quadratic function $\eta x^{T}M^{T}Mx$. Such an assumption is commonly encountered in the literature addressing the stabilization of nonlinear systems using LMIs [33].
Consider
$Q^{B}_{P}(x,u):=$
$\max\left\{h(x),[u^{T}\;\bar{v}^{T}(x)]\mathbf{D}^{\dagger T}\mathbf{x}^{T}_{1,1,k}P^{\!-\!1}\mathbf{x}_{1,1,k}\mathbf{D}^{\dagger}\left[\begin{gathered}u\\ \bar{v}(x)\end{gathered}\right]\!-\!1\right\},$ (26)
where $P\in\mathbb{R}^{n_{x}\times n_{x}}$ is a positive-definite matrix. The consideration of the parameterization (4.4) is inspired from Proposition 1 and (21), in which we specify $h(x)=\max_{i}\{H_{x,i}x-h_{x,i}\}$ so that $\{x\in\mathbb{R}^{n_{x}}|h(x)\leq 0\}$ is identical to $\{x\in\mathbb{R}^{n_{x}}|H_{x}x\leq h_{x}\}$. The second argument of the max function indicates that we firstly consider a quadratic CBF candidate $x^{T}P^{-1}x-1$, and then use the open-loop representation (21) and the result in Proposition 1 to get an SACBF candidate.
Now, we consider the following convex optimization problem:
$\min_{P,R_{1},G_{2}}$ $\;-\log\operatorname{det}(P)$ (27a)
$\mathrm{s.t.}\;$ $\left[\begin{array}[]{cccc}P&0&R^{T}_{1}\mathbf{x}^{T}_{1,1,k}&PM^{T}\\ &I_{n_{v}}&G^{T}_{2}\mathbf{x}^{T}_{1,1,k}&0\\ &*&P&0\\ &*&*&I_{n_{x}}/\eta\end{array}\right]\succeq 0$ (27f)
$\left[\begin{array}[]{cccc}h^{2}_{u,i}P&0&R^{T}_{1}\mathbf{u}^{T}_{0,1,k}H^{T}_{u,i}&PM^{T}\\ &I_{n_{v}}&G^{T}_{2}\mathbf{u}^{T}_{0,1,k}H^{T}_{u,i}&0\\ &*&1&0\\ &*&*&I_{n_{x}}/\eta\end{array}\right]\succeq 0$ (27k)
$\quad\quad i=1,...,\dim(h_{u})$ (27l)
$\left[\begin{array}[]{cc}h^{2}_{x,i}&H_{x,i}P\\ &P\end{array}\right]\succeq 0,\;i=1,...,\dim(h_{x})$ (27o)
$\left[\begin{array}[]{c}\mathbf{x}_{0,1,k}\\ \mathbf{v}_{0,1,k}\end{array}\right]R_{1}=\left[\begin{array}[]{c}P\\ 0\end{array}\right]$ (27t)
$\left[\begin{array}[]{c}\mathbf{x}_{0,1,k}\\ \mathbf{v}_{0,1,k}\end{array}\right]G_{2}=\left[\begin{array}[]{c}0\\ I_{n_{v}}\end{array}\right]$ (27y)
$P\succeq 0,\;R_{1}\in\mathbb{R}^{k\times n_{x}},\;G_{2}\in\mathbb{R}^{k\times n_{v}},$ (27z)
where “*” indicates that the lower triangular part of the matrix is the symmetric counterpart of the upper triangular part and $H_{x,i},\;h_{x,i}$ refer to the $i$th row of $H_{x},\;h_{x}$. Based on the optimization problem (27), we can construct an SACBF $Q^{B}_{P}$, as stated in the next result.
Proposition 4.
Consider the system (19), the SACBF parameterization (4.4), and the optimization problem (27). Suppose that $\mathrm{rank}(\mathbf{D})=n_{x}+n_{u}+n_{v}$ and Assumptions 3-4 hold. Then, problem (27) is feasible, and its optimal solution $P^{*}$ makes $Q^{B}_{P^{*}}$ an SACBF with the corresponding safe set $\{x\in\mathbb{R}^{n_{x}}|x^{T}P^{*-1}x\leq 1\}$.
Proof.
The proof is based on the standard argument of using LMIs to synthesize CBFs [23]. By letting $R_{1}=G_{1}P$ and $K=\mathbf{u}_{0,1,k}G$, (27t) and (27y) are equivalent to (23).
We consider the condition
$x^{+T}P^{-1}x^{+}\leq 1,\;\forall x^{T}P^{-1}x\leq 1,\;\forall v^{T}(x)v(x)\leq\eta x^{T}M^{T}Mx,$ (28)
where $x^{+}$ is the successor state of the closed-loop system (22). By sequentially using the S-procedure [34], utilizing the Schur complement [34], multiplying both sides by $\mathrm{diag}([P\;I_{n_{v}}\;I_{n_{x}}])$, and applying the Schur complement once more, we get that (27f) is sufficient to ensure (28). The condition (28) further implies the invariance of the safe set for the closed-loop system (22).
Then, by a similar argument we can prove that (27l) is a sufficient condition for
$H_{u}K\bar{v}(x)\leq h_{u},\;\forall x\in\{x\in\mathbb{R}^{n_{x}}|x^{T}P^{-1}x\leq 1\}$ (29)
to hold. The condition (29) means that the controller $u=K\bar{v}(x)$ satisfies the input constraint for all states in the safe set. Besides, (27o) is equivalent to the condition
$\{x\in\mathbb{R}^{n_{x}}|x^{T}P^{-1}x\leq 1\}\subseteq X.$ (30)
The above arguments prove that $B:=x^{T}P^{*-1}x-1$ is a CBF. Meanwhile, by comparing the right-side of (4.4) and the form of $B$, and noting that the open loop system (19) can be represented by (21), we have the relation $Q^{B}_{P^{*}}(x,u)=\max\{h(x),\;B(f(x,u))$. Finally, according to Proposition 1, we can conclude that $Q^{B}_{P^{*}}$ is an SACBF.
We now prove the feasibility of (27). The feasibility of (27t) and (27y) follows from (23). Regarding the other constraints, since (i) the linearized system of (19) is controllable, (ii) $\lim_{x\rightarrow 0}\frac{||v(x)||_{2}}{||x||_{2}}=0$ from Assumption 4, and (iii) the origin is within the interior of $X$, we can always replace $P$ by $\alpha_{P}P$ with $\alpha_{P}\in(0,1)$ to make the conditions (28)-(30) satisfied. This proves the feasibility of (27f)-(27o). ∎
5 Performance analysis and guarantee under learning errors
In the previous section, we have introduced three different learning-based methods for computing an SACBF. However, learning errors are unavoidable and can arise from a combination of factors including insufficient data, loss function mismatch, and suboptimal optimization. These errors have the potential to invalidate the learned SACBF. To address this problem, we propose a systematic analysis to evaluate the robust safety performance of (3) when it includes an inaccurate approximation $Q^{B}_{\omega}$. This analysis further leads to a practical approach for handling learning errors through state constraint tightening followed by SACBF constraint relaxation.
Our analysis is inspired by the concept of input-to-state safety (ISSf) [35], which was originally developed for studying set invariance in the presence of disturbances. Before introducing the framework, we point out that the analysis of safety performance can be applied to the SL method in Section 4.1 and the expert-guided learning method in Section 4.2, while the RL method in Section 4.3 is excluded. We will explain the reason for this exclusion at the end of this section.
First, we relax the constraint in the original optimization-based controller (3), based on the intuition that it may no longer be possible to render (3) recursively feasible in the presence of learning errors. Define a control policy with a relaxed SACBF constraint as
$\pi^{\mathrm{r}}_{\theta,\omega}(x):=\arg\min_{u\in U}$ $\;Q_{\theta}(x,u)$
$\mathrm{s.t.}\;$ $Q^{B}_{\omega}(x,u)\leq\kappa(\varepsilon),$ (31)
where $\kappa(\cdot)$ is a $\mathcal{K}_{\infty}$ function and $\varepsilon\geq 0$. The notation $\varepsilon$ will be used to quantify the learning error, and $\kappa$ will be elucidated later.
Then, similar to how CBFs are extended to ISSf CBFs [35], we introduce error-to-state safety SACBFs (ESSf SACBFs), capturing the set invariance when $Q^{B}_{\omega}$ is different from $Q^{B}$.
Definition 3 ($\varepsilon$-ESSf SACBF).
A function $\hat{Q}^{B}(\cdot,\cdot):\mathcal{X}\times U\to\mathbb{R}$ is called an $\varepsilon$-error-to-state safe SACBF ($\varepsilon$-ESSf SACBF) with a corresponding safe set $\hat{\mathcal{S}}_{Q}$, if there exists a $\mathcal{K}_{\infty}$ function $\kappa(\cdot)$ such that the pair $(\hat{Q}^{B},\;\hat{\mathcal{S}}_{Q})$ satisfies the following conditions:
(i) $\hat{\mathcal{S}}_{Q}$ is non-empty.
(ii) $\min_{u\in U}\hat{Q}^{B}(x,u)\leq\kappa(\varepsilon),\;\forall x\in\hat{\mathcal{S}}_{Q}$.
(iii) For any $x\in\hat{\mathcal{S}}_{Q}$, any $u\in U$ satisfying $\hat{Q}^{B}(x,u)\leq\kappa(\varepsilon)$ ensures that $f(x,u)\in\hat{\mathcal{S}}_{Q}$.
Now, we are ready to give our main results on the ESSf property of $Q^{B}_{\omega}$ learned by the two methods presented in Sections 4.1 and 4.2.
Theorem 1 (ESSf for the SL method).
Under Assumption 1, consider the controlled system $x_{t+1}=f(x_{t},\pi^{\mathrm{r}}_{\theta,\omega}(x_{t}))$, an SACBF $Q^{B}$, which is induced by a CBF $B$ from (8) with $\beta_{B}<1$, and an approximation $Q^{B}_{\omega}$ of $Q^{B}$. Suppose that
$|Q^{B}(x,u)-Q^{B}_{\omega}(x,u)|\leq\varepsilon,\;\forall(x,u)\in\mathcal{X}\times U$ (32)
holds. Letting $\kappa(\varepsilon)=\frac{1+\beta_{B}}{1-\beta_{B}}\varepsilon$, we have:
(i) The approximation $Q^{B}_{\omega}$ is an $\varepsilon$-ESSf SACBF. The corresponding safe set is $\mathcal{S}_{\omega}=\{x\in\mathcal{X}|B(x)\leq\frac{2\beta_{B}}{1-\beta_{B}}\varepsilon\}$.
(ii) The optimization-based controller (5) is recursively feasible with the initial condition $x\in\mathcal{S}_{\omega}$, i.e., $\mathcal{S}_{\omega}$ is forward invariant for the controlled system $x_{t+1}=f(x_{t},\pi^{\mathrm{r}}_{\theta,\omega}(x_{t}))$.
(iii) Furthermore, if $B$ is the CBF for (1) under the tightened state constraint $x\in X_{\varepsilon}:=\{x\in\mathbb{R}^{n}|h(x)+\frac{2\beta_{B}}{1-\beta_{B}}\varepsilon\leq 0\}$, the controlled system satisfies the original constraints $x_{t}\in X$ and $u_{t}\in U$, $\forall t\in\mathbb{N}$.
Proof.
According to Lemma 1, $\mathcal{S}_{Q}=\mathcal{S}_{B}$. The non-emptiness of $\mathcal{S}_{\omega}$ follows directly from the non-emptiness of $\mathcal{S}_{B}$. Then, we will prove the feasibility of (5) for any $x\in\mathcal{S}_{\omega}$ (Condition (ii) of Definition 3). From (8) and using the properties of CBFs, we have $\min_{u\in U}Q(x,u)\leq\max\{h(x),\;B(x)\}\leq B(x),\;\forall x\in\mathcal{X}$. For all $x\in\mathcal{S}_{\omega}$, it follows from (32) that
$\min_{u\in U}Q^{B}_{\omega}(x,u)$ $\leq\min_{u\in U}Q^{B}(x,u)+\varepsilon$
$\leq B(x)+\varepsilon\leq\underbrace{\frac{{1+{\beta_{B}}}}{{1-{\beta_{B}}}}\varepsilon}_{\kappa(\varepsilon)}.$
Next, let $x\in\mathcal{S}_{\omega}$ and $u$ be any input satisfying $u\in U$ and $Q^{B}_{\omega}(x,u)\leq\kappa(\varepsilon)$. We have
$B(f(x,u))\leq\beta_{B}Q^{B}(x,u)\leq\beta_{B}Q^{B}_{\omega}(x,u)+\beta_{B}\varepsilon\leq\frac{2\beta_{B}}{1-\beta_{B}}\varepsilon.$
The above inequalities prove the forward invariance of the set $\mathcal{S}_{\omega}$ and the satisfaction of Condition (iii) of Definition 3. The above arguments prove the statements (i) and (ii).
Furthermore, if $B$ is a CBF under the tightened constraint $x\in X_{\varepsilon}$, we derive that $B(x)-\frac{2\beta_{B}}{1-\beta_{B}}\varepsilon\geq h(x)$, which further implies that the safe set $\mathcal{S}_{\omega}$ of the learned SACBF $Q^{B}_{\omega}$ is contained in the original state constraint set $X$. As $\mathcal{S}_{\omega}$ is forward invariant, the infinite-time safety of the controlled system follows. ∎
Theorem 2 (ESSf for the expert-guided learning method).
Under Assumption 2, consider the controlled system $x_{t+1}=f(x_{t},\pi^{\mathrm{r}}_{\theta,\omega}(x_{t}))$ and approximations $Q^{B}_{\omega}$ and $q_{\omega}$ of $Q^{B}$ and $q$ in (9). Suppose that
$Q^{B}_{\omega}(x,\pi_{\mathrm{s}}(x))\leq\beta q_{\omega}(x)+\varepsilon,\;\forall x\in\mathcal{S}_{0}$ (33a)
$q_{\omega}(f(x,u))\leq Q^{B}_{\omega}(x,u)+\varepsilon,\;\forall(x,u)\in\mathcal{S}_{0}\times U$ (33b)
$q_{\omega}(x)\geq\frac{2}{1-\beta}\varepsilon,\;\forall x\in\mathcal{X}\setminus\mathcal{S}_{0}$ (33c)
holds. Letting $\kappa(\varepsilon)=\frac{1+\beta}{1-\beta}\varepsilon$, we have:
(i) The approximation $Q^{B}_{\omega}$ is an $\varepsilon$-ESSf SACBF. The corresponding safe set is $\mathcal{S}_{\omega}=\{x\in\mathbb{R}^{n}|q_{\omega}(x)\leq\frac{2}{1-\beta}\varepsilon\}$.
(ii) The optimization-based controller (5) is recursively feasible with the initial condition $x\in\mathcal{S}_{\omega}$, i.e., $\mathcal{S}_{\omega}$ is forward invariant for the controlled system $x_{t+1}=f(x_{t},\pi^{\mathrm{r}}_{\theta,\omega}(x_{t}))$.
(iii) Furthermore, if $h(x)+\frac{2}{1-\beta}\varepsilon\leq q_{\omega}(x),\;\forall x\in\mathcal{X}$, the controlled system satisfies the original constraints $x_{t}\in X$ and $u_{t}\in U$, $\forall t\in\mathbb{N}$.
Proof.
(33c) leads to $\mathcal{S}_{\omega}\subseteq\mathcal{S}_{0}$. Consider any $x\in\mathcal{S}_{\omega}$, we have
$\min_{u\in U}Q_{\omega}^{B}(x,u)\leq Q_{\omega}^{B}(x,\pi_{\mathrm{s}}(x))\leq\beta q_{\omega}(x)+\varepsilon\leq\underbrace{\frac{1+\beta}{1-\beta}\varepsilon}_{\kappa(\varepsilon)},$
which proves the feasibility of (5).
Moreover, consider any $x\in\mathcal{S}_{\omega}$ and any $u\in U$ satisfying $Q^{B}_{\omega}(x,u)\leq\kappa(\varepsilon)$. We have
$q_{\omega}(f(x,u))\leq Q^{B}_{\omega}(x,u)+\varepsilon\leq\frac{2}{1-\beta}\varepsilon.$
The rest of the proof can be completed using the same reasoning as that of the proof of Theorem 1. ∎
The condition in (32) quantifies the local approximation quality of the regression model $Q^{B}_{\omega}$, which is often assumed in SL literature [36, 37, 23]. The two conditions in (33a) and (33b) are motivated by (9c) and (9d), where we assume that the constraints in (9c) and (9d) are violated and the degree of violation is limited by $\varepsilon$. Besides, the condition (33c) is also justifiable, as it is introduced to ensure that the safe set of $Q^{B}_{\omega}$ is a subset of $\mathcal{S}_{0}$.
Remark 4.
Different from existing work that primarily addresses the robustness and ISSf for learning-based control under external disturbances [5], input disturbances [37], or inaccurate observations [38], our findings offer a systematic way for designers to guarantee the safety of the relaxed safety-filtered policy $\pi^{\mathrm{r}}_{\theta,\omega}$ when there are perturbations on the safety constraints, by appropriately tightening the state constraint. More importantly, we explicitly quantify the degrees of tightening and relaxation, based on the parameters $\beta$ and $\beta_{B}$ of the given CBF and the safe policy $\pi_{\mathrm{s}}$ as well as the bound $\varepsilon$ on the learning error. In particular, to ensure safety for the SL method, the state constraint must be tightened to $h(x)+\frac{2\beta_{B}}{1-\beta_{B}}\varepsilon\leq 0$, while the safety constraint in the safety filter should be relaxed by adjusting $Q^{B}_{\omega}(x,u)\leftarrow Q^{B}_{\omega}(x,u)-\frac{1+\beta_{B}}{1-\beta_{B}}\varepsilon$. Similarly, in the expert-guided learning method, the state constraint should be tightened to $h(x)+\frac{2}{1-\beta}\varepsilon\leq 0$, and (9b) should be adjusted to (33c), with the relaxed SACBF $Q^{B}_{\omega}(x,u)\leftarrow Q^{B}_{\omega}(x,u)-\frac{1+\beta}{1-\beta}\varepsilon$ in the safety filter.
Remark 5.
The proposed constraint tightening approach is inspired by robust MPC [37], which ensures both recursive feasibility and safety. In contrast, our optimization-based control framework enforces safety through an SACBF constraint, rather than imposing state constraints over a finite prediction horizon. As a result, recursive feasibility alone does not guarantee safety, highlighting the need for constraint relaxation. Broadly speaking, state constraint tightening contributes to safety, whereas SACBF constraint relaxation addresses feasibility concerns.
At the end of this section, we explain why the RL method cannot be included in the proposed ESSf framework. To make $Q^{B*}$ in (16) an SACBF, the optimal control problem (4.3) should be un-discounted, which leads to a non-contractive Bellman equation (17) [39]. As a consequence, the learning error of $Q^{B*}$ could become unbounded. Besides, in Theorem 1 we require $\beta_{B}<1$ while in the reachability formulation (4.3), $\beta_{B}=1$. The above two factors make the ESSf analysis inapplicable to the RL method.
6 Refining policies with SACBF constraints
Given a valid SACBF obtained by the methods in the previous section, in this section, we update $\theta$ to refine the feasible policy $\pi_{\theta,\omega}$, bringing it closer to the optimal policy that solves (3).
A convenient approach for designing $Q_{\theta}$, allowing $\theta$ to be updated using most existing unconstrained policy-based RL algorithms is to consider a Euclidean distance objective function $Q_{\theta}(x,u):=||u-\pi_{\theta}(x)||_{2}$, where $\pi_{\theta}:\mathcal{X}\rightarrow\mathcal{U}$ is an explicit controller (e.g., a neural network controller) with the parameter $\theta$ [7]. This approach, however, could significantly compromise the optimality of the projected policy $\pi_{\theta,\omega}$. A less conservative approach is to make $Q_{\theta}$ approximate the constrained optimal value function $Q^{*}$ defined in (6), so that the policy derived from (3) more closely approximates the optimal constrained policy. To this end, we present a unified approach to transform unconstrained value-based RL algorithms into constrained ones, utilizing the obtained SACBF.
Normally, to obtain $Q^{*}$, as is shown in [40, Chapter 7.4.1], one typically needs to recursively compute the backward reachable sets $X_{k},\;k=0,1,...$ and enforce the constraint $f(x,u)\in X_{k}$ during the Bellman iteration. This procedure is in general intractable for nonlinear systems. Instead, we replace the state constraints $h(x_{t})\leq 0,\;\forall t\in\mathbb{N}$ in (3) with the obtained SACBF constraints $Q_{\omega}^{B}(x_{t},\pi(x_{t}))\leq 0,\;\forall t\in\mathbb{N}$. Note that this will be an equivalent transformation if $Q_{\omega}^{B}=Q^{B*}$. The benefit of this transformation is that the constraints during the Bellman iteration become fixed to $Q_{\omega}^{B}(x_{t},\pi(x_{t}))\leq 0$. This is a result of the inherent invariance of the safe set of $Q_{\omega}^{B}$.
By combining the following expression of $\bar{J}^{*}$:
$\bar{J}^{*}\left(x\right):=\inf_{\pi}J^{\pi}\left(x\right)\;\mathrm{s.t.}\;Q_{\omega}^{B}(x_{t},\pi(x_{t}))\leq 0,\pi(x_{t})\in U,t\in\mathbb{N},$ (34)
and the expression (3) of $\pi_{\theta,\omega}$, we know that $\pi_{\theta,\omega}$ is optimal for the (34) if $Q_{\theta}$ equals $\bar{Q}^{*}(x,u):=g(x,u)+\gamma\bar{J}^{*}(f(x,u))$.
The constrained optimal state-action value function $\bar{Q}^{*}$ satisfies the following constrained Bellman equation:
$\bar{Q}^{*}(x,u)=\Gamma\bar{Q}^{*}:=g(x,u)+\gamma$ $\min_{u^{+}\in U}\bar{Q}^{*}(f(x,u),u^{+})$
$\mathrm{s.t.}\;Q_{\omega}^{B}(f(x,u),u^{+})\leq 0,$ (35)
where $\Gamma$ is the Bellman operator. It is easy to show that $\Gamma$ is a monotonous contraction mapping. As a consequence, the uniqueness of the solution to (6) holds and the constrained Bellman iteration $\bar{Q}_{k+1}=\Gamma\bar{Q}_{k}$ converges to $\bar{Q}^{*}$ for any real-valued and bounded $\bar{Q}_{0}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ [39, Proposition 2.1.1].
With this property, we can update $\theta$ to minimize the average Bellman residual over $\mathcal{S}_{\omega}\times U$. Formally speaking, we find
$\theta^{*}:=\arg\min_{\theta}\int_{\mathcal{S}_{\omega}\times U}(q_{\theta}(x,u)$ $-Q_{\theta}(x,u))^{2}dxdu,$ (36a)
where $\text{where }q_{\theta}(x,u)=g(x,u)+\gamma$ $\min_{u^{+}\in U}Q_{\theta}\left(f(x,u),u^{+}\right)$
[…]
Problem (36a) is bi-level and intractable to solve exactly. In practice, following standard offline value iteration algorithms [22], we iteratively update $\theta$. In each iteration, $q_{\theta}$ is treated as the target value, and $\theta$ is updated such that the difference between $q_{\theta}$ and $Q_{\theta}$ is minimized. This ultimately forms the proposed constrained fitted Q iteration (constrained FQI) algorithm (Algorithm 1). Furthermore, just as standard FQI can be adapted to constrained FQI, other online value-based RL algorithms, such as constrained deep Q-learning, constrained approximate SARSA, and Lagrangian RL, can likewise be designed accordingly.
Algorithm 1 Constrained fitted Q iteration
1:Given the SACBF $Q^{B}_{\omega}$, the state and input constraint sets $X$ and $U$, $\mathcal{X}$, the sample set $\mathcal{D}\subseteq\mathcal{S}_{\omega}\times U\times\mathcal{S}_{\omega}$, the learning rate $\zeta>0$, the error threshold $\varepsilon_{\mathrm{QI}}>0$, and the maximum number of updates $K\in\mathbb{N}^{+}$
2:Initialize $\theta_{0}$, $k\leftarrow 0$
3:Repeat at each iteration $k$ $q_{s}\leftarrow\Gamma Q_{\theta_{k}}(x_{s},u_{s})$ for each $(x_{s},u_{s},x^{+}_{s})\in\mathcal{D}$ $\theta_{k+1}\leftarrow\arg\min_{\theta}\sum_{s=1}^{|\mathcal{D}|}\left(q_{s}-Q_{\theta_{k}}\left(x_{s},u_{s}\right)\right)^{2}$ $k\leftarrow k+1$
4:Until $|Q_{\theta_{k+1}}(x,u)-Q_{\theta_{k}}(x,u)|\leq\varepsilon_{\mathrm{QI}},\;\forall(x,u)\in\mathcal{X}\times U$, or $k>K$
5:Output $Q_{\theta_{k}}$
7 Comparison with other approaches and limitations
Comparison with learning CBFs: Machine learning has recently been developed for synthesizing CBFs (safety filters) with or without model uncertainties. A common methodology to synthesizing permissive CBFs is to link CBFs with value functions of optimal control problems, such as the approaches in [8, 12, 5, 17] and the current paper. The value function can then be approximated by RL or SL. Some authors assume the knowledge of a nominal model and a valid CBF for that model [14, 13]. Then, RL is used to learn the discrepancies between the CBF constraint associated with the nominal model and that related to the real model. However, all the above methods can only enable the design of model-based safety filters. This limitation is due to the inherent nature of CBFs, which need the model information to enforce the invariance condition. In contrast, in the current paper we fundamentally propose a new direct data-driven safe control framework, in which an SACBF is employed to evaluate the feasibility of input signals without using model information.
The paper in [20] uses discriminating hyperplanes (a series of linear constraints) to represent safe constraints in safety filters. This approach eliminates the dependence on any specific safety certificate, and consequently, on the model. The hyperplanes can be updated by RL, in which the reward is designed such that unsafe policies are highly penalized. However, the RL method of [20] can still be understood as a reward shaping method for managing safety, which lacks formal guarantees regarding constraint satisfaction. Besides, approximating nonlinear constraints by linear constraints may lead to a rather conservative policy. In contrast, the proposed optimization-based control framework, which involves a nonlinear program, provides a formally sound and more general method to enhance safety and to eliminate dependence on an explicit model.
Comparison with integrating RL and MPC: The combination of MPC and RL can have various kinds of forms. [41] uses RL algorithms to update a parameterized nonlinear MPC scheme. It is shown in [41] that the parameterized MPC scheme can produce safe and stabilizing policies as long as the RL algorithm updates parameters in a personally-defined safe and stable set. Such a set, however, is usually problem-dependent. It is possible to use LMIs to determine such a set for linear systems. Different from [41], which parameterizes all terms of MPC including the terminal cost, the prediction model, and the constraints, [36] only parameterizes the terminal cost and uses approximate value iteration to learn it offline. Similarly, [42] adopts approximate policy iteration to learn the terminal cost offline. It is proven in [36, 42] that the resulting MPC controller makes the system safe and asymptotically stable if the approximation error is bounded and the MPC horizon is large enough. The above combinations, however, fail to solve the online computational problem faced by MPC, since their policies are still determined by online optimization over a long prediction horizon. In comparison, the proposed optimization-based control approach yields significant online computational benefits by using two state-action value functions $Q_{\theta}$ and $Q^{B}_{\omega}$ to approximate the value function and safety constraints implicitly defined by the parameterized MPC scheme.
Moreover, we acknowledge that, unlike parametrized MPC, which employs an explicit prediction model to derive optimal and safe policies for long-term goals, our optimization-based control framework lacks explainability in the resulting policy. However, our method offers greater flexibility, as it allows the integration of any RL algorithm into the control synthesis.
Comparison with safe RL: Existing methods for safe RL include using reward shaping [25, 7], Lagrangian methods combined with policy gradient or actor-critic algorithms [43], interior-point optimization [44], and safety filtering [18, 45]. Additionally, a small body of work explores direct stochastic optimization of neural network controllers over finite horizons, improving safety at sampled states but demanding significant computational resources [46]. However, as discussed in the introduction, because learning algorithms operate stochastically, learning-based control—particularly when an explicit model is unavailable—requires safety filters to regulate policy execution and to ensure reliable safety. The proposed optimization-based control framework is compatible with all the safe RL methods mentioned above, and provides formal safety assurance under learning errors.
Comparison with our previous work [23]: In [23], the definition of SACBFs is proposed, which contributes to a safety enhancement framework analogous to (3). The current work expands [23] in the following two aspects. First, all the learning-based methods for SACBFs proposed in Section 4 totally remove the assumption on the availability of a nominal model, which is required in [23]. Second, in [23], the robustness to learning errors is addressed by a new CBF with a contractive safe set, which needs a stronger invariance condition than invariant sets. In contrast, the current work establishes an ESSf analysis framework using the fundamental robustness property of the SACBF, resulting in significantly less conservative constraint tightening than [23].
Limitations: One downside is that safety constraints will inevitably be violated when collecting samples of state transitions to train the controller. This may be inappropriate for real-world applications where maintaining safety is essential throughout the learning process. Using simulators to generate sample data during learning can solve this issue, although the gap from sim to real needs further robustness analysis.
We extend the barrier certificate, originally used to characterize state safety, to the SACBF, which captures both state-input safety. However, this extension introduces a drawback: it necessitates sampling and learning in the state-input space rather than just the state space, thereby increasing computational and sample complexity.
8 Case study
In this section, we verify the proposed data-driven approaches for synthesizing safety filters and their application to safe learning-based control for an autonomous vehicle moving in a 2-D space containing obstacles.
8.1 Model
We consider the kinematic vehicle model [5]:
$\dot{p}_{x}$ $=v\cos(\Psi)$ (37)
$\dot{p}_{y}$ $=v\sin(\Psi)$
$\dot{v}$ $=a$
$\dot{\Psi}$ $=v\tan(\delta)/L,$
where $L=0.1$. The state vector $x$ includes the position $p_{x}$, $p_{y}$, the speed $v$, and the yaw angle $\Psi$. The acceleration $a$ and the steering angle $\delta$ are the inputs. The input constraints are given by $-5\leq a\leq 2$ and $|\delta|\leq\pi/4$. The state constraints are specified by $|p_{x}|\leq 2.6$, $|p_{y}|\leq 2.6$, $0\leq v\leq 1$, and $|\Psi|\leq\pi$, as well as the requirement of avoiding some obstacles shown in Fig. 2. According to the considered state constraints, the function $h$ is as follows:
[…]
$\left.\max_{i=1,2,3,4}\{r^{2}_{i}-(p_{x}-c_{x,i})^{2}-(p_{y}-c_{y,i})^{2}\}\right\}.$
Here, $r_{i}$ denotes the radius of each obstacle, while $c_{x,i}$ and $c_{y,i}$ represent the $x$- and $y$-coordinates of the center of each obstacle. The state constraint is tightened to $\{x|h(x)+0.2\leq 0\}$ when learning SACBFs. The target set is $X_{\mathrm{tar}}=\{x\in\mathbb{R}^{4}|p^{2}_{x}+p^{2}_{y}\leq 0.1^{2}\}$. The system is discretized using the Euler method with sampling time 0.05 s.
8.2 Synthesizing SACBFs
The synthesis of the SACBF is performed using three approaches, including the expert-guided learning approach presented in Section 4.2, the RL approach described in Section 4.3, and the LMI approach of Section 4.4. The SL approach in Section 4.1 is not considered because it is unrealistic to assume knowing a CBF without knowing the model in this example. For the expert-guided approach, we adopt artificial potential fields (APF) [47] to design the expert controller. The APF-based controller has a certain capability for obstacle avoidance but is prone to falling into local optima, which can either reside within obstacle regions or be located far from the target.
To get the training data sets $\mathcal{S}_{\mathrm{s}}$, $\mathcal{X}_{\mathrm{s}}$ for the expert-guided approach, and $U_{\mathrm{s}}$, we start the system from 10000 randomly generated initial conditions inside $\mathcal{X}=\{x\in\mathbb{R}^{4}\mid\;|p_{x}|\leq 3,\;|p_{y}|\leq 3,\;-0.2\leq v\leq 1.2,\;|\Psi|\leq\pi\}$ and get the trajectories over 200 time steps. For the remaining learning models, including using RL to learn the SACBF $Q^{B}$, learning the optimal value function $Q$, learning the reference control policies, and identifying the model in the subsequent statistical tests, we use uniformly random sampling to get $10^{6}$ state-input samples from $\mathcal{X}$.
[…]
In Figs. 2(a)-(d), we illustrate the neural SACBFs obtained by the expert-guided learning approach and the RL approach, and their corresponding safe sets. To visualize the 4D sets, we display their shapes on the $p_{x}$-$p_{y}$ plane with fixed $v=0.5$ and $\Phi=0$. Both safe sets avoid intersections with obstacles and wall areas. The safe set derived from the expert controller is smaller than that learned through the RL approach. This difference arises because the RL method theoretically approximates the maximal safe set, whereas the expert controller (APF) is not always feasible within the maximal safe set, which is illustrated by the red curves in Fig. 2(a).
(a) The boundaries (blue curves) of the safe set $\{x|q_{\omega}(x)\leq 2\varepsilon/(1-\beta)\}$ learned from the expert controller. The green curves represent the safe trajectories of the APF controller, while the red curves and points represent the unsafe trajectories and initial states.
(b) The value of $q_{\omega}$ learned from the expert controller.
(c) The boundaries (blue curves) of the safe set $\{x|\min_{u\in U}Q^{B}_{\omega}(x,u)\leq 0\}$ learned by RL
(d) The value of $\min_{u\in U}Q^{B}_{\omega}(\cdot,u)$ learned by RL.
Figure 2: The SACBFs and their corresponding safe sets in the $p_{x}$-$p_{y}$ plane with $v=0.5$ and $\Psi=0$. The green area represents the obstacles and walls. In the contour figures, values below zero mean that the positions are feasible.
Figure 3: The boundaries (blue curves) of the safe set $\{x|x^{T}P^{-1}x\leq 1\}$ of the quadratic SACBF (4.4) obtained by solving LMIs
For the LMI approach, since Assumption 3 cannot be satisfied for the original vehicle model in (37), we make some simplifications. In particular, (i) only the subsystems of $p_{y}$ and $\Psi$ are considered, (ii) the speed is fixed to a constant $0.5$, (iii) the obstacles are removed, and (iv) the approximation $\tan\delta\approx\delta$ is used. Then, the resulting simplified system is of the form (19), with the nonlinear term $\sin\Psi$. After using the input signal $\delta(t)=0.1\sin t$ to excite the system, we get the state-input data $\mathbf{D}$ (defined in Lemma 2) with the time horizon $20$. The persistency of excitation condition is satisfied. By solving the LMI problem (27), we obtain $Q^{B}_{P}$ with $P=\left[\begin{array}[]{cc}6.25&-0.2996\\ -0.2996&0.1176\end{array}\right]$. The corresponding safe set is visualized in Fig. 3. As observed, the LMI approach is conservative, as it restricts the SACBF candidate to quadratic forms.
8.3 Closed-loop simulation
Hereafter, we evaluate the performance of the proposed optimization-based controller (3). We construct two reference control policies obtained using: (i) APF and (ii) deep deterministic policy gradient (DDPG) [50]. The APF contains two elements: attractive potential fields, which guide the vehicle toward the target set, and repulsive potential fields, which prevent the vehicle from hitting the obstacles. To highlight the performance of the safety filter, we only apply the attractive potential field to get the reference controller (referred to as unsafe APF). Similarly, in the case of DDPG, we do not incorporate safety constraints during the training process (referred to as unsafe DDPG). For DDPG, the stage cost is designed as $g(x,u)=\max\{0,p^{2}_{x}+p^{2}_{y}-0.1^{2}\}$, with a discount factor $\gamma=0.99$. For comparison, we evaluate the filtered policies against: (i) the APF method incorporating both attractive and repulsive potential fields (referred to as the safe APF); (ii) the DDPG policy trained with a penalty on constraint violations in the cost function (referred to as safe DDPG) [51].
Fig. 4(a) demonstrates the performance of the 3DSF learned using the expert-guided learning approach. When applying the unsafe policy, the trajectories exhibit some constraint violations. In contrast, the learned SACBF enables the vehicle to smoothly avoid obstacles, except for the initial condition $p_{x}=2,\;p_{y}=0.5$ (the rightmost black dot). To explain this constraint violation, we note that the expert controller (safe APF) also exhibits constraint violations when starting from this initial state. Consequently, the 3DSF, which is learned based on this expert controller, recognizes this initial state as unsafe and therefore fails to refine the reference policy.
In comparison, the 3DSF learned by RL refines the reference policy for all the initial states. The trajectory starting from the rightmost initial state in Fig. 4(b) demonstrates the superiority of the proposed 3DSF learned by RL. In particular, the safe APF policy makes the trajectory fail into a local optimum inside the right obstacle, while the unsafe APF policy with the 3DSF circumvents the local optimum and steers the system to the target.
In Figs. 4(c-d), we show the trajectories of the vehicle controlled by the DDPG policies. It is found that although for some initial states safe DDPG successfully plans a safe trajectory, some trajectories fail to converge to the target set. This illustrates that it is hard to balance the task performance and constraint satisfaction by naively adding penalties to the stage cost during training. In contrast, the safety filter significantly enhances the safety performance of the unsafe DDPG policy while ensuring that the policy reaches the target.
(a) Closed-loop trajectories. Dashed blue curve: unsafe APF; Solid blue curve: unsafe APF with 3DSF learned from the expert controller; Red curve: safe APF; Black dots: initial states.
(b) Closed-loop trajectories. Dashed blue curve: unsafe APF; Solid blue curve: unsafe APF with 3DSF learned by RL; Red curve: safe APF; Black dots: initial states.
(c) Closed-loop trajectories. Dashed blue curve: unsafe DDPG; Solid blue curve: unsafe DDPG with 3DSF learned from the expert controller; Red curve: safe DDPG; Black dots: initial states.
(d) Closed-loop trajectories. Dashed blue curve: unsafe DDPG; Solid blue curve: unsafe DDPG with 3DSF learned by RL; Red curve: safe DDPG; Black dots: initial states.
Figure 4: Comparison of closed-loop vehicle trajectories under different controllers and different safety filters.
8.4 Statistical evaluation and comparison with indirect data-driven safety filters
(a) Trajectories of the vehicle controlled by the unsafe APF policy filtered by the 3DSF.
(b) Trajectories of the vehicle controlled by the unsafe APF policy filtered by the indirect safety filter.
Figure 5: Comparison of closed-loop vehicle trajectories under the 3DSF and the indirect safety filter.
Finally, we statistically compare the proposed 3DSF against an indirect data-driven safety filter. In this subsection, the SACBF is learned by the RL approach because it has been illustrated in the previous subsection that the SACBF learned by the RL approach results in a larger safe set. To design the indirect data-driven safety filter, we intend to learn a standard CBF based on an approximate model. We first use a neural network with [128 128 128] “elu” hidden layers to identify the kinematic model. After 30 epochs of training, we get an approximate model with a sufficiently small ($2.38\cdot 10^{-4}$) mean square error. Then, following the approaches of [12, 8, 23], the approximate model is used as a prediction model in the reachability problem [^4] The reachability problem (4.3) is defined over an infinite horizon. To make it solvable for each sampled state, we truncate the horizon at 20. While a longer horizon provides a better approximation of the maximal safe set, a horizon of 20 balances between sample complexity and safety performance. (4.3). We solve (4.3) with the initial state $x_{0}$ being each state sample obtained in Section VIII.B. Then, we get the target value for the neural CBF, which is parameterized by a neural network with [128 128 32] “tansig” hidden layers.
Table 1: Comparison of safety rate, total cost, and average online CPU time for different control policies and safety filters. The total cost is defined as the sum of $p^{2}_{x}+p^{2}_{y}$ over 1000 time steps. The safety rate is defined as the ratio between the number of successful trajectory plans (without constraint violations) and the number of all trajectory plans. “SF” means “safety filter”.
Controller Performance No SF Indirect SF 3DSF
Unsafe APF
Safety Rate
Total Cost
CPU Time
58.16%
114.30
0.0043 ms
79.43%
136.59
1.60 ms
100%
154.25
2.19 ms
Safe APF
Safety Rate
Total Cost
CPU Time
81.56%
373.39
0.0060 ms
80.85%
521.46
1.51 ms
100%
155.87
2.21 ms
Unsafe DDPG
Safety Rate
Total Cost
CPU Time
70.92%
120.33
0.39 ms
79.43 %
141.22
2.21 ms
96.45%
187.54
2.49 ms
Safe DDPG
Safety Rate
Total Cost
CPU Time
82.27%
560.60
0.36 ms
89.36%
571.45
2.49 ms
100%
611.90
2.61 ms
Constrained FQI
Safety Rate
Total Cost
CPU Time
100%
173.85
4.95 ms
We randomly sample 141 initial states that lie in the intersection of the safe set of the SACBF and the safe set of the CBF. Fig. 5 shows the trajectories controlled by the unsafe APF controller with the 3DSF and the indirect counterpart, respectively. Notably, the results demonstrate the absence of undesired equilibria, limit cycles, or unbounded trajectories.
As has been mentioned in Section 2, (3) can be utilized either as a safety filter (3DSF) for a pre-obtained reference control policy or as a policy generator to determine suboptimal control inputs greedy to $Q_{\theta}$. Therefore, we apply the constrained FQI (Algorithm 1) to get the approximation $Q_{\theta}$ of the constrained optimal value function $\bar{Q}^{*}$ defined in (6). This makes (2.3) a greedy policy optimization problem.
The results in Table 1 provide a detailed comparison of the safety rate, total cost, and average CPU time for those trajectories starting from the sampled initial states. The total cost reflects the ability of governing the vehicle to the target set. The key findings are summarized below:
  • Both the 3DSF and the indirect data-driven safety filter can in general reduce the rate of constraint violation while not significantly degrading the performance regarding the total cost. When combined with safety filters, even unsafe policies (e.g., Unsafe APF and Unsafe DDPG) achieve safety rates comparable to their safe counterparts.
  • Most importantly, the 3DSF significantly improves the safety rate compared to the indirect counterpart. The worse performance of the indirect counterpart is likely due to the superposition effect of the model error and the CBF learning error.
  • Constrained FQI achieves high safety rates (100%) and a lower total cost (173.85) compared to safe and unsafe DDPG with the 3DSF. This validates that including the SACBF constraint in the training process of RL could improve the task performance.
9 Conclusions and future work
We have proposed an optimization-based control framework that contains an SACBF constraint to ensure safety for learning-based control methods. The main advantages lie in its universal applicability to RL and SL methods for designing controllers, as well as its online computational efficiency. Three learning algorithms have been developed within the optimization-based control framework, making the controller strive for optimal performance while ensuring safety to the greatest extent possible. We have analyzed the theoretical properties regarding the robustness of SACBF and translated the results to error-to-state safety (ESSf) of the proposed control framework w.r.t. learning errors. Simulations conducted on an obstacle avoidance problem demonstrate the aforementioned advantages.
Future work will focus on (i) increasing the scalability to higher dimensional systems, (ii) examining the effect of external uncertainties, and (iii) designing distributed optimization-based control for multi-agent systems subject to unknown dynamics and nonlinear constraints.